A farmer has a rectangular garden plot surrounded by 200 ft of fence. Find the lenght and with of the gfarden if its area is 2400ft(square)
Answer by checkley71(8403)

2X+2Y=200 & XY=2400 THEN 2X=200-2Y OR X=(200-2Y)/2 OR X=100-Y
SUBSTITUTE 100-Y FOR X IN THE SECOND EQUATION & WE GET (100-Y)Y=2400 OR
100Y-Y~2=2400 OR Y~2-100Y+2400 OR (Y-60)(Y-40)=0 OR Y=60 & Y=40 THE
X*60=2400 OR X=40 WHEN Y=60 & X*40=2400 OR X=60 WH3EN Y=40 THUS THE DIMENTIONS
OF THE PLOT ARE 40 BY 60 FEET.